OpenStudy (anonymous):

PLEAS HELP QUICK!!!!!! MEDAL for correct answer A yearbook printer charges based on the number of pages printed. Here is a table that shows the cost of some recent yearbooks. Number of pages 50 100 150 200 Cost 1.55 2.25 2.95 3.65 Which equation gives the cost, y, in terms of the number of pages, x? A.y = 0.014x + 49.15 B.y = 0.014x + 1.55 C.y = 0.014x + 0.85 D.y = 58.8x + 0.85

OpenStudy (anonymous):

Alright, I'm in the mood to use the process of elimination.

OpenStudy (anonymous):

Alright let's look at the cost for 50 pages. We can eliminate the answer B because if we use any number besides 0, we would end up with a number bigger than 1.55. We can eliminate the answer A because we're adding 49.15 to any response. That means that the number of pages would have to be negative which is possible in my own universe (Let's not talk about that) We can eliminate the answer D because we're multiplying 50 by 58.8 which is already surpassing 1.55. This leaves the answer to be C. Please sue me if I'm wrong :)

OpenStudy (anonymous):

ok let me get my lawyer

OpenStudy (anonymous):

just kidding let me check

OpenStudy (anonymous):

CORRECT thank you will you help me with 4 more

OpenStudy (anonymous):

@*Shadow*

OpenStudy (anonymous):

Sure :)

OpenStudy (anonymous):

Which is an equation in point-slope form for the given point and slope? Point: ; (8,-4) Slope:-5/6 A.y+4=-5/6(x+8) B.y-4=-5/6(x+8) C.y+4=-5/6(x-8) D.y-4=5/6(x-8)

OpenStudy (anonymous):

@*Shadow*

OpenStudy (anonymous):

Sorry, was checking something.

OpenStudy (anonymous):

What is the point-slope form of the equation of a line that passes through the point (1, 4) and has a slope of -3? A.y-1=-3(x-4) B.y+4=3(x+1) C.y+4=-3(x+1) D.y-4=-3(x-1) so what do you think for this 2?

OpenStudy (anonymous):

@*Shadow* you still here?

OpenStudy (abdullah1995):

@NunChuckBoy3101 you still want help ?

OpenStudy (anonymous):

yes please would you like a medal to?

OpenStudy (abdullah1995):

i don't care about medals. only like to help others :)

OpenStudy (anonymous):

ok

OpenStudy (abdullah1995):

So first job is to expand all the options you have. multiply it all out.

OpenStudy (anonymous):

because i can open another question so you and @*Shadow* can both have them

OpenStudy (anonymous):

Which is an equation in point-slope form for the given point and slope? Point: ; (8,-4) Slope:-5/6 A.y+4=-5/6(x+8) B.y-4=-5/6(x+8) C.y+4=-5/6(x-8) D.y-4=5/6(x-8) so it would be D

OpenStudy (anonymous):

@abdullah1995

OpenStudy (abdullah1995):

can you wait 10-15 minutes ? i have to go really important. but i will be back

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

what do you think @*Shadow*

OpenStudy (abdullah1995):

@NunChuckBoy3101 you still here ?

OpenStudy (anonymous):

hey i am back now

OpenStudy (anonymous):

@abdullah1995

OpenStudy (abdullah1995):

ok

OpenStudy (abdullah1995):

so first thing you want to do is to expand every option you have. expand the brackets

OpenStudy (anonymous):

ok i already did those will you help me with this 1

OpenStudy (abdullah1995):

ok post the question

OpenStudy (anonymous):

A yearbook printer charges based on the number of pages printed. Here is a table that shows the cost of some recent yearbooks: number of pages 50 100 150 200 cost 1.25 1.85 2.45 3.05 Which equation gives the cost, y, in terms of the number of pages, x? A.y = 0.012x + 49.985 B.y = 0.012x + 1.25 C.y = 83.3x + 0.65 D.y = 0.012x + 0.65

OpenStudy (anonymous):

@abdullah1995

OpenStudy (abdullah1995):

ok lets try to solve this. consider the number of pages your x value and charges as y value to make co-ordinates that is (50, 1.25) ( 100, 1.85) and so on use only two pairs so i choose 50 1.25 and 100 and 1.85 now take the slope

OpenStudy (abdullah1995):

the options that are given are actually in the form y=mx+b ; once you find the slope you will find the m value. looking at the options it is either going to be 0.012 or 83.3

OpenStudy (anonymous):

so b

OpenStudy (abdullah1995):

nope

OpenStudy (abdullah1995):

what did you get as your slope value ?

OpenStudy (anonymous):

84

OpenStudy (anonymous):

hang on i subtracted wrong

OpenStudy (abdullah1995):

its not 84. try again

OpenStudy (abdullah1995):

(y2-y1)/(x2 - x1) = m

OpenStudy (abdullah1995):

(50, 1.25) ( 100, 1.85) (1.85-1.25)/(100-50) = m . if you solve it you will get 0.012

OpenStudy (anonymous):

i know i just did i have no idea ho i got 84 lol

OpenStudy (anonymous):

how*

OpenStudy (anonymous):

ok i got 0.012

OpenStudy (abdullah1995):

if you are still confused ask the question. solving for marks is not enough. otherwise we can move ahead

OpenStudy (anonymous):

no we can move ahead

OpenStudy (abdullah1995):

ok great ! now we have m value. we also have (x, y). there are a lot of pairs you can just take one of them. y = mx+b ( input x, y and m and solve for b )

OpenStudy (anonymous):

?

OpenStudy (abdullah1995):

(1.85) = (100)(0.012) + b (1.85)-(100*0.012) = b b = 0.65

OpenStudy (anonymous):

oh wait its y = 0.012x + 0.65

OpenStudy (abdullah1995):

did you get it ? more importantly did you understand the whole process ?

OpenStudy (anonymous):

yes let me submit

OpenStudy (abdullah1995):

alright sweet.

OpenStudy (anonymous):

yes it was correct thank you if i have any more questions like this i will tag you in it

OpenStudy (abdullah1995):

sure thing. if i dont see the tag send me a private message and we can work through it

OpenStudy (anonymous):

thank you im about to tag you in something

OpenStudy (abdullah1995):

sure thing